Cramers Rule: 5. Use Cramers Rule to find x,y and z for the following system of equations. X 2 7x + 2y - z= -1 ។ 6x + 5y +
Answer #1

4x+ 23-7 51 117. Gu+5g+2 = 46 - Sk-4943 1 = -5 ㅋ 2. G 5 -54 3 11-2) : +1 ① ( 15 + 4 ) + 2 (1645) - 4 (-24 +25) 3M9+ 46-멋 9 45PAGE NO DATE 101 Dr 11 2 5 -5 -4 (1) ( 1544) -1 (4645) + (-4) (404-25) 19- 206 73 -198 고 Dy = IC - 3 4. ( 46 +5) - (-) (1645)